# Towards Quantum Accelerated Large-scale Topology Optimization

### v1

@@ 5 Conclusion
We have presented a new methodology capable of efficiently solving TO problems while enabling the integration of quantum computing to harness potential quantum advantages. This work has targeted on addressing large-scale, multi-material TO challenges for 3D continuum structures, representing a significant departure from previous research that leveraged quantum computing in TO [34, 35, 46, 21, 36]. Those studies primarily tackled discrete truss structures or linear or small-scale continuum problems, overlooked the additional cost induced by constructing QUBO, and exclusively dealt with single-material designs. At the heart of this new methodology is the MDW decomposition, which effectively addresses the escalating computational cost associated with using classical MILP solvers to directly solve the master problems in the DVTO-MT framework [20] that features a multi-cut formulation and adaptive trust regions and can significantly reduce iteration counts, and in turn, the time required for solving the governing PDE via FEM. The MDW decomposition breaks down each MILP problem into $n_{D}$ local sub-problems (BIP) and one global sub-problem (LP). The BIP problems, which are the most computationally intensive, can be accelerated by quantum computing through solving their equivalent QUBO formulations.
@@
We have evaluated the computational efficiency and solution quality of our method by designing 3D bridges using both single and multiple materials. As it directly handles binary design variables, our method naturally produces clear-cut 0/1 topologies, unlike SIMP [19], the predominant TO approach, which relaxes binary variables into continuous ones (varying between 0 and 1) and requires post-processing, such as Heaviside projection [47], to revert to binary representations. In its classical implementation, our method achieves comparable objective function values while reducing overall computation time by at least an order of magnitude compared to SIMP [19]. When benchmarked against DVTO-MT [20] on the time required for solving the master problems, the present method can achieve speedups from one to three orders of magnitude. Remarkably, the MDW decomposition maintains low computational time even for the largest problem tested ($150\times 600\times 150$, yielding 54 million design variables), where the MILP solver fails to return a solution. The proposed quantum implementation has the potential to further accelerate solving the resultant BIP sub-problems by about another order of magnitude. Due to current constraints in accessing real quantum hardware, we verified the validity of the formulated QUBO using QAOA [45]. All the speedups noted above consistently become more pronounced with increasing problem size and complexity (moving from single-material to multi-material designs). This suggests that the new method and quantum computing will play an increasingly valuable role in addressing the scale and complexity of real-world TO applications.
@@
Another important advantage of the proposed MDW decomposition is that each resulting BIP (local sub-problem) involves only local constraints. As a result, the corresponding QUBO formulation features a sparse Hamiltonian, eliminating the need for all-to-all qubit connectivity, a key consideration for near-term quantum hardware. Additionally, each local sub-problem can be solved independently within each MDW iteration, allowing all QUBOs to be processed in parallel.
@@
It is also important to note that our estimates of quantum speedup include the time required to construct the QUBO problem, $T_{\text{QUBO}}$, i.e., the time to generate the QUBO’s coefficient matrix $\mathbf{Q}$. Since our method allows $\mathbf{Q}$ to be sparse, $T_{\text{QUBO}}$ should theoretically scale linearly with the problem size. However, practical code implementation can affect the actual scalability of $T_{\text{QUBO}}$. To achieve the desired linear scalability for $T_{\text{QUBO}}$, an improved code implementation is necessary, as discussed in §4.3. With this improvement, we anticipate an even more substantial quantum speedup.
In this work, we have focused on reducing both the total number of iterations and the time required to solve the optimization sub-problems at each iteration. By decreasing the iteration count, we have effectively reduced the number of repeated PDE solves, thereby lowering the overall time spent on PDE solving. Further improvements in efficiency are expected if the per-iteration PDE solving can also be accelerated. To this end, integrating quantum-accelerated PDE solvers [48, 49, 50] into TO represents a promising direction for future work, offering the potential to further enhance the efficiency and scalability of TO.

### v2

@@ 5 Conclusion
Two key factors largely determine the total computational time in TO: (1) the number of iterations required to achieve convergence, which dictates the number of PDE solver (e.g., FEM) calls, and (2) the time needed to solve the optimization problem at each iteration. The DVTO-MT framework developed in our prior work [20] significantly reduces the number of iterations, and consequently, the time spent solving the governing PDE via FEM. Building upon this framework, the MDW decomposition proposed in the present work further substantially reduces the time needed to solve the optimization problem at each iteration. Accordingly, a major contribution of this work is the development of an efficient TO methodology that markedly enhances computational efficiency by reducing both the iteration count and per-iteration optimization time. This work has targeted on addressing large-scale, multi-material TO challenges for 3D continuum structures.
Another key contribution lies in enabling the integration of quantum computing to leverage potential quantum advantages. The MDW decomposition partitions each MILP problem into $n_{D}$ local sub-problems (BIPs) and one global sub-problem (LP). The BIP sub-problems, being the most computationally intensive, can be accelerated via quantum computing by solving their equivalent QUBO formulations. The formulated QUBOs are compatible with both quantum annealing and gate-based quantum computing. This represents a significant advancement over previous efforts that applied quantum computing to TO [34, 35, 49, 50, 51, 52, 21, 36, 53, 54], which primarily focused on discrete truss or small-scale continuum problems and were limited to single-material designs.
@@
We have evaluated the computational efficiency and solution quality of our method by designing 3D bridges using both single and multiple materials. As it directly handles binary design variables, our method naturally produces clear-cut 0/1 topologies, unlike SIMP [19], the predominant TO approach, which relaxes binary variables into continuous ones (varying between 0 and 1) and requires post-processing, such as Heaviside projection [55], to revert to binary representations. In addition, our method achieves comparable objective function values while reducing overall computation time by at least one order of magnitude compared to SIMP [19]. When benchmarked against the original DVTO-MT [20] on the time required for solving the master problems, the present method with the MDW decomposition can achieve speedups from one to three orders of magnitude. And it maintains low computational time even for the largest problem tested ($150\times 600\times 150$, yielding 54 million design variables), for which the classical MILP solver fails to return a solution. The proposed quantum implementation has the potential to further accelerate solving the resultant BIP sub-problems by about another order of magnitude. Due to current constraints in accessing real quantum hardware, we have verified the validity of the formulated QUBO using QAOA [47]. All the speedups noted above consistently become more pronounced with increasing problem size and complexity (moving from single-material to multi-material designs). This suggests that the new method and quantum computing will play an increasingly valuable role in addressing the scale and complexity of real-world TO applications.
@@
Another important advantage of the proposed MDW decomposition is that each resulting BIP (local sub-problem) involves only local constraints. As a result, the corresponding QUBO formulation features a sparse Hamiltonian, eliminating the need for all-to-all qubit connectivity, a key consideration for near-term quantum hardware. In this regard, the proposed approach surpasses previous studies, including our earlier work [21] and others [36], which produced QUBOs with dense Hamiltonians requiring all-to-all connectivity and incurring quadratically scaled construction costs with respect to the number of design variables. Moreover, since all local sub-problems are independent within each MDW iteration, their QUBOs can be processed in parallel, further enhancing computational efficiency.
@@
It is important to note that our estimates of quantum speedup account for the time required to construct the QUBO problem, $T_{\text{QUBO}}$, i.e., the time needed to generate the QUBO coefficient matrix $\mathbf{Q}$. Because our method produces a sparse $\mathbf{Q}$, $T_{\text{QUBO}}$ should theoretically scale linearly with problem size. However, in practice, implementation details can influence the actual scalability. Achieving this ideal linear scaling will require an improved code implementation, as discussed in §4.3, which we plan to pursue next. With this enhancement, we expect an even greater quantum speedup.
Finally, we have focused on reducing both the total number of iterations and the per-iteration optimization time. The reduction in iteration count effectively decreases the number of repeated PDE solves, thereby lowering the overall computational cost. Further efficiency gains could be realized by accelerating the PDE solving itself. In this regard, integrating quantum-accelerated PDE solvers [56, 57, 58] into TO presents a promising avenue for future research, offering the potential to further enhance the efficiency and scalability of TO. Along different lines, integrating with variational quantum algorithms [54], which enable single-loop parallel searches for optimal configurations that satisfy physical constraints through quantum entanglement, offers another potential future direction. In addition, quantum machine learning (QML) approaches could be leveraged to achieve further acceleration, such as the framework proposed by Sukulthanasorn and Terada [59]. Their approach employs quantum neural networks based on parameterized quantum circuits to learn the complex relationships between the objective function, sensitivity, and design variables, can scale to arbitrary problem sizes with a fixed number of qubits, and has been shown to substantially accelerate optimality criteria (OC)–based TO.
CRediT authorship contribution statement
Zisheng Ye: Writing – original draft, Conceptualization, Methodology, Software, Investigation, Validation, Formal analysis, Data curation. Wenxiao Pan: Writing – original draft, Writing – review & editing, Supervision, Conceptualization, Methodology, Investigation, Funding acquisition, Resources, Project administration.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgment
Data availability
The codes and examples that support the findings of this study are available in the link: https://github.com/Pan-Group-UW-Madison/QCTO-MDW.
Appendix A Prescribed Tolerance in Eq. (30)
In all numerical results presented in this paper, the tolerance in Eq. (30), which governs the termination of the MDW iterations, was set to $10^{-6}$. To assess the sensitivity of the outer DVTO–MT iteration (the $k$-iteration) to the tolerance prescribed for the inner MDW loop (the $l$-loop), we conducted additional tests with tolerance values ranging from $10^{-3}$ to $10^{-7}$. For illustration, we considered the single-material design by MDW-CC, as described in §4.2, with the discretization resolution of $100\times 400\times 100$. The results are summarized in Table A.1. As the tolerance is relaxed, only a slight increase in the total iteration count ($N_{\text{Iter}}$) or the resulting objective function value ($f$) is observed. These results demonstrate that the proposed method is robust with respect to approximations in the Lagrange multipliers.
Table A.1: Sensitivity of the objective function value ($f$) and total iteration count ($N_{\text{Iter}}$) to the tolerance prescribed in Eq. (30).
Tolerance $N_{\text{ite}}$ $f$
$10^{-7}$ 41 13.571
$10^{-6}$ 47 13.624
$10^{-5}$ 67 13.552
$10^{-3}$ 53 13.948
Remarks: We note that the results reported for the tolerance value $10^{-6}$ differ slightly from those presented in Table 2. This discrepancy is primarily due to that the Gurobi optimizer (v11.0.0) [44] employs stochastic heuristics to accelerate the search for optimal solutions, which prevents exact reproducibility across independent runs. In addition, this set of tests were conducted on a workstation equipped with two Intel® Xeon® CPU E5-2698 v4 processors, using 32 MPI processes per run. Variations in domain decomposition associated with the number of processing cores may also introduce minor differences in the results.